# Toward Exploring the Code Understanding Capabilities of Pre-trained Code Generation Models

### v1

@@ Abstract
Recently, large code generation models trained in a self-supervised manner on extensive unlabeled programming language data have achieved remarkable success. While these models acquire vast amounts of code knowledge, they perform poorly on code understanding tasks, such as code search and clone detection, as they are specifically trained for generation. Pre-training a larger encoder-only architecture model from scratch on massive code data can improve understanding performance. However, this approach is costly and time-consuming, making it suboptimal. In this paper, we pioneer the transfer of knowledge from pre-trained code generation models to code understanding tasks, significantly reducing training costs. We examine effective strategies for enabling decoder-only models to acquire robust code representations. Furthermore, we introduce CL4D, a contrastive learning method designed to enhance the representation capabilities of decoder-only models. Comprehensive experiments demonstrate that our approach achieves state-of-the-art performance in understanding tasks such as code search and clone detection. Our analysis shows that our method effectively reduces the distance between semantically identical samples in the representation space. These findings suggest the potential for unifying code understanding and generation tasks using a decoder-only structured model.
Toward Exploring the Code Understanding Capabilities of Pre-trained Code Generation Models
Jiayi Lin1, Yutao Xie1, Yue Yu2, Yibiao Yang3, Lei Zhang1, 1International Digital Economy Academy, 2National University of Defense Technology, 3Nanjing University, Correspondence: jiayilin1024@gmail.com
1 Introduction
@@
Figure 1: Timeline of code-large models. For visual clarity, only the largest size of representative models is plotted. The orange arrows indicate the development trends of encoder-only code understanding models, while the green arrows represent the trends of decoder-only code generation models.
Recently, there has been an explosion of pretrained code generation models that have achieved remarkable success Nijkamp et al. (2023); Guo et al. (2024); Lozhkov et al. (2024). These models are pre-trained in a self-supervised manner on vast amounts of code data, exhibiting robust code generation capabilities. They have been developed into tools that assist software developers in their daily programming tasks, such as GitHub Copilot.
Although these decoder-only code generation models have acquired substantial code knowledge during pre-training, they perform poorly on code comprehension tasks, as they are specifically trained for generation tasks using autoregression. Some studies have shown that better code understanding performance can be achieved by pre-training a larger encoder-only model from scratch with more training data Hoffmann et al. (2022); Lin et al. (2024). However, this process is not only expensive but also time-consuming, making it suboptimal. As shown in Figure 1, numerous increasingly larger code generation models have been proposed, such as the 1.1B SantaCoder Allal et al. (2023), the 15B StarCoder Li et al. (2023), and the 70B CodeLlama Rozière et al. (2023). These models are significantly larger than the current largest code understanding models, namely the 1.3B CodeSage Zhang et al. (2024) and the 1.5B CoLSBERT Lin et al. (2024), and have been exposed to more code data. However, leveraging these pre-trained code generation models for code understanding tasks has not been previously explored.
@@
In this paper, we investigate the effective application of a pre-trained code generation model with a decoder-only Transformer architecture to code understanding tasks. As illustrated in Figure 2, we explore two methods for extracting code representations from the decoder-only Transformer: 1) using the encoding of the last token after passing through the decoder, and 2) using the average of the encodings of all tokens after passing through the decoder. Additionally, we introduce CL4D, which employs contrastive learning to continue pre-training code generation models, thereby enhancing their representational capabilities. We follow Gao et al. Gao et al. (2021) by selecting code snippets that are semantically different but close to the query in the representation space as hard negatives to improve training efficiency.
@@
To ensure objectivity, we compare code understanding models with code generation models of equivalent size on two understanding tasks: code search and clone detection. Experimental results indicate that our proposed CL4D method effectively utilizes decoder-only code generation pre-trained models for code understanding tasks. After continued pre-training on a small code corpus, these models outperform pre-trained encoder-only code understanding models on most tasks. Furthermore, we observe that code understanding tasks can benefit from larger decoder-only models.
@@
  • We identify the value of pre-trained code generation models and are the first to effectively apply them to code understanding tasks.
  • Our proposed CL4D method improves the representation ability of decoder-only models, enabling them to outperform encoder-only models in code understanding tasks without fine-tuning.
@@
  • Our findings demonstrate the potential for a decoder-only model to unify both code understanding and code generation tasks.
  • We will release model training code and pre-trained model parameters to contribute to the further research of code understanding large models.
2 Related works
Code Represention for Programming Languages Recently, obtaining programming language representations for a wide range of downstream code understanding tasks has garnered significant attention. Models such as CodeBERT Feng et al. (2020), StarEncoder Li et al. (2023), CodeSage Zhang et al. (2024), and CoLSBERT Lin et al. (2024) have been inspired by the success of the Masked Language Modeling (MLM) objective in natural language processing tasks. They optimize this objective on serialized code data, treating the code as text. Some studies have introduced additional tasks commonly used in text understanding, such as the replaced token detection objective Devlin et al. (2019) or the next sentence prediction objective Clark et al. (2020), for source code. However, recognizing that code is different from text, some works propose leveraging the structural information of code during training to enhance semantic learning. For instance, GraphCodeBERT Guo et al. (2021) uses data flow to encode variable value flows in functions, TreeBERT Jiang et al. (2021) extracts syntactic structure from the abstract syntax tree (AST) to guide the training process, and SynCoBERT Wang et al. (2021) and UniXcoder Guo et al. (2022) incorporate the serialized AST as part of the input for model training.
While these works based on encoder-only Transformer contribute to the development of code understanding tasks, they are limited by the “small” model sizes and the relatively “small” code corpora they are trained on, typically consisting of millions of samples. The largest model among them is only 1.5B parameters, namely CoLSBERT. By contrast, decoder-only code generation models, such as the 15B parameter StarCoder Li et al. (2023) trained on a 1T token corpus, have significantly larger model sizes and training data. Motivated by these limitations, our research aims to transfer the extensive code knowledge from pre-trained code generation models to code understanding tasks. We utilize their decoder-only architecture to obtain code representations for downstream code understanding tasks, achieving state-of-the-art performance.
(a) Last
(b) Avg
Figure 2: Different ways to get code representation in decoder-only models.
Contrastive Learning for Code Contrastive learning is widely employed in representation learning of deep learning networks. Recent works have applied contrastive learning to code corpora, proposing various methods to construct positive or negative samples to enhance code representation. For example, CoSQA Huang et al. (2021) constructs positive samples by rewriting the query, SynCoBERT Wang et al. (2021) and Code-MVP Wang et al. (2022) create positive sample pairs across different code modalities using diverse information obtained during program compilation, and UniXcoder Guo et al. (2022), inspired by SimCLR Chen et al. (2020), uses dropout to construct positive samples in the latent vector space. Additionally, ChatDANCE Wang et al. (2023) uses large language models such as ChatGPT to synthesize positive code samples. CodeRetriever Li et al. (2022), R2 Dong et al. (2023), and CodeSage Zhang et al. (2024) construct hard negatives for queries using rankers. In this paper, to address the insufficient representation ability of decoder-only models due to the unidirectional attention mechanism, we select code snippets with different semantics but close to the query in the representation space as hard negatives.
3 Methodology
@@
In this section, we explore obtaining encoding from decoder-only pre-trained code generation models to address code understanding tasks, with models ranging from 125 million to several billion parameters. We first introduce methods to effectively obtain code representation from decoder-only models, followed by how to continue pre-training these models using contrastive learning to enhance their representation capabilities.
3.1 Model Architecture
As illustrated in Figure 2, we investigated two methods for obtaining code representation from decoder-only models: 1) using the embedding of the last token, and 2) averaging the embeddings of all tokens. Encoder-only models like BERT typically use the embedding of the first token $[CLS]$ in the last hidden layer as the feature representation of the input code snippet, utilizing a bidirectional attention mechanism. By contrast, decoder-only models use a unidirectional attention mechanism, where only subsequent tokens can refer to previous information. Consequently, only the last token aggregates the information of the entire sample. In the first method, we use the embedding of the last token in the last hidden layer as the feature representation of the input code snippet. In the second method, we calculate the average of the embeddings of all input tokens in the last hidden layer as the feature representation.
Figure 3: Architecture of the dual-encoder model.
Following previous work Guo et al. (2021); Zhang et al. (2024), we adopt a dual-encoder architecture to continue pre-training the code encoding model. As depicted in Figure 3, this architecture consists of two Transformer decoder modules with shared weights to encode the input code snippet. In our experiments, we initialize the Transformer decoder module with the pre-trained code generation models.
3.2 Training
The traditional left-to-right training method for predicting the next token limits the representation capability of code generation models. To fully leverage the extensive code knowledge in the existing open-source code generation models and to avoid the high costs of retraining code understanding models, we employ contrastive learning to further train these models, thereby improving their representation abilities.
3.2.1 Training data
@@
We train decoder-only Transformer models on six programming languages from The Stack dataset Kocetkov et al. (2022): Python, Java, Go, PHP, JavaScript, and Ruby. To enhance the ability of model to differentiate between natural language and programming language, we use Tree-Sitter [^1] https://tree-sitter.github.io/tree-sitter/ to extract bimodal data from The Stack dataset for training, i.e., $(query,code)$ pairs. The query is the first line of the function docstring, which approximates the query of user. Additionally, we apply the following filtering rules to improve the quality of the training corpus:
  • Exclude samples where the query is not in English.
  • Exclude samples where the query is less than 3 tokens or more than 256 tokens.
  • Exclude samples where the function name contains “test”.
  • Exclude constructor samples.
  • Exclude samples where the function body is fewer than 3 lines.
Given that pre-trained code generation models have already learned a vast amount of code knowledge, we construct millions of data samples for continued pre-training. The statistical results of the pre-training data are shown in the Table 1.
Language #Functions
Python 714,352
Java 718,902
Go 715,664
Php 707,424
JavaScript 707,778
Ruby 64,157
Total 3,628,277
Table 1: Statistics of training data.
3.2.2 Contrastive Learning
@@
Contrastive learning is the primary method for enhancing the representation ability of models Zhang et al. (2023). Here, we employ contrastive learning tasks to train pre-trained code generation models, addressing the limited representation ability of decoder-only Transformer due to the unidirectional attention mechanism.
@@
Specifically, we randomly sample n samples from the training corpus to form a batch of data $B={(q_{1},c_{1}),...,(q_{n},c_{n})}$. For a $(q_{i},c_{i})$ pair, we take $c_{i}$ as the positive sample $c^{+}$ for $q_{i}$, and the other $n-1$ codes in the batch, excluding $c_{i}$, as negative samples for $q_{i}$. Additionally, following Gao et al. Gao et al. (2021), we construct hard negative samples $c^{h}$ for $q_{i}$, i.e., code snippets that are close to the query in the representation space but have different semantics from the query. The loss function can be formulated as:
@@ 4 Experiments
  • Q1: What is the optimal method for extracting code representations from decoder-only Transformer?
  • Q2: How well do raw pre-trained code generation models perform on code understanding tasks?
  • Q3: To what extent does contrastive learning enhance the code representations of decoder-only Transformer?
  • Q4: Why does our approach work?
4.1 Experimental Setup
Understanding Tasks We evaluate the code understanding capabilities of decoder-only Transformer on two primary tasks:
  • Code Search: This task aims to find the most semantically similar code snippet in a candidate codebase based on a given natural language query. We perform experiments on two datasets: CodeSearchNet (CSN) Husain et al. (2019) and CoSQA Huang et al. (2021). CSN is a multi-language code corpus collected from GitHub, including Python, Java, Go, PHP, JavaScript, and Ruby. CoSQA contains real user queries from Microsoft Bing search logs, with code snippets from CSN. We use Mean Reciprocal Rank (MRR) as the evaluation metric, which is widely adopted in previous studies Guo et al. (2022); Zhang et al. (2024).
  • Clone Detection: This task aims to identify the most semantically similar code snippet in a candidate codebase based on a given code snippet as query. We conduct experiments on the POJ-104 dataset Mou et al. (2016), designed to measure semantic similarity between code snippets. Mean Average Precision (MAP) is the evaluation metric for this task.
Baseline We compare the code understanding capabilities of encoder-only models and decoder-only models of similar sizes, and evaluate the improvements in understanding capabilities provided by the proposed CL4D method for decoder-only models.
@@ 4.1 Experimental Setup
  • CodeBERT Feng et al. (2020) (125M) is trained with masked language modeling and replaced token detection on six programming languages in CodeSearchNet.
  • GraphCodeBERT Guo et al. (2021) (125M) is trained with additional code structure understanding tasks, including edge prediction and node alignment.
  • UniXcoder Guo et al. (2022) (125M) is trained with multiple language modeling and contrastive learning tasks. It is an encoder-decoder architecture model, where only the encoder is used for code understanding tasks via a special token $<encoder\_only>$.
  • CodeSage Zhang et al. (2024) (1.3B) is trained with masked language modeling and contrastive learning on the extensive code corpus from The Stack dataset.
@@
  • CodeGPT Lu et al. (2021) (125M) is trained on CodeSearchNet with the same structure and training objectives as GPT-2 Radford et al. (2019).
  • CodeGen Nijkamp et al. (2023) (350M) is a model for program analysis trained on extensive natural language and programming language data. We choose the 350M version to facilitate comparison with similarly sized encoder-only models.
  • SantaCoder Allal et al. (2023) (1.1B) is trained with multi-query attention and fill-in-the-middle on Java, JavaScript, and Python subsets of The Stack, with additional near-deduplication and comments-to-code filtering.
  • phi-1 Gunasekar et al. (2023) (1.3B) is a code generation model trained using textbook data sourced from the internet and textbooks synthesized by GPT-3.5.
The network architectures of these models are detailed in Table 3.
Model size L. A. H.
Encoder-only CodeBERT 125M 12 12 768
GraphCodeBERT 125M 12 12 768
UniXcoder 125M 12 12 768
CodeSage 1.3B 24 16 2048
Decoder-only CodeGPT 125M 12 12 768
CodeGen 350M 20 16 1024
SantaCoder 1.1B 24 16 2048
phi-1 1.3B 24 32 2048
Table 3: Model Architecture.
Training details During the pre-training phase, we utilize the tokenizer from each decoder-only model directly for tokenization. These models are trained using the AdamW optimizer, with a learning rate set to 2e-5. We train the decoder-only models for 2 epochs on the dataset presented in Table 1, employing a batch size of 64. The experiments are conducted on 8 GPUs (80G A100). The longest pre-training duration is observed for phi-1, which takes approximately 3 days. In the fine-tuning phase, we adhere to the experimental settings of UniXcoder to minimize the impact of hyper-parameters on the fine-tuning results.
4.2 Q1: What is the optimal method for extracting code representations from decoder-only Transformer?
Decoder-only models generate the next token in an autoregressive manner, using a unidirectional attention mechanism that references only previous information during encoding. To determine the optimal method for obtaining code representations in decoder-only Transformer models, we investigated two approaches: decoder-only last token (Last) and decoder-only average (Avg), as depicted in Figure 2. Additionally, we examined the impact of padding on the effectiveness of code representations.

### v2

# Towards Better Code Understanding in Decoder-Only Models with Contrastive Learning

@@ Abstract
Recent advances in large-scale code generation models have led to remarkable progress in producing high-quality code. These models are trained in a self-supervised manner on extensive unlabeled code corpora using a decoder-only architecture. However, despite their generative strength, decoder-only models often exhibit limited performance on code understanding tasks such as code search and clone detection, primarily due to their generation-oriented training objectives. While training large encoder-only models from scratch on massive code datasets can improve understanding ability but remains computationally expensive and time-consuming. In this paper, we explore a more efficient alternative by transferring knowledge from pre-trained decoder-only code generation models to code understanding tasks. We investigate how decoder-only architectures can be effectively adapted to learn discriminative and semantically meaningful code representations. To this end, we propose CL4D, a contrastive learning framework tailored to strengthen the representation capabilities of decoder-only models. Extensive experiments on multiple benchmark datasets demonstrate that CL4D achieves competitive or superior performance compared to existing methods on representative code understanding tasks, including code search and clone detection. Further analysis reveals that CL4D substantially improves the semantic alignment of code representations by reducing the distance between semantically similar code snippets. These findings highlight the feasibility of leveraging decoder-only models as a unified backbone for both code generation and understanding.
Codehttps://github.com/JiayiLin1024/CL4D
Introduction
Recent years have witnessed a surge of large-scale pre-trained code generation models. These models are trained in a self-supervised manner on vast code corpora using a decoder-only architecture and have demonstrated remarkable success (Nijkamp et al. 2023; Guo et al. 2024; Lozhkov et al. 2024). They have been widely adopted in developer-assistance tools, such as GitHub Copilot [^1] https://github.com/features/copilot and Cursor [^2] https://cursor.com/, to support everyday programming tasks.
@@ Introduction
Figure 1: Timeline of code-large models. For visual clarity, only the largest size of representative models is plotted. The orange arrow indicate the development trends of encoder-only code understanding models, while the green arrow represent the trends of decoder-only code generation models.
Despite their outstanding generative performance, decoder-only models often struggle with code understanding tasks, including code search and code clone detection. This limitation arises from their auto-regressive training objectives, which primarily emphasize next-token prediction rather than semantic comprehension. Previous studies have shown that stronger performance on code understanding tasks can be achieved by pre-training encoder-only models from scratch on large-scale code datasets (Hoffmann et al. 2022; Lin et al. 2024). However, such training is computationally expensive and time-consuming, making it impractical for widespread adoption. As illustrated in Figure 1, the landscape of code models has evolved rapidly: from 1.1B SantaCoder (Allal et al. 2023) and the 15B StarCoder (Li et al. 2023) to the 70B CodeLlama (Rozière et al. 2023). These models are significantly larger and more data-rich than the current leading code understanding models, such as the 1.3B CodeSage (Zhang et al. 2024) and the 1.5B CoLSBERT (Lin et al. 2024). Nevertheless, how to effectively leverage these powerful pre-trained decoder-only code generation models for code understanding tasks remains an open and unexplored problem.
@@
In this paper, we investigate the adaption of pre-trained decoder-only code generation models to code understanding tasks. We explore two methods for extracting code representations from the decoder-only Transformer: 1) using the encoding of the last token, and 2) using the average encodings of all tokens. With these code representations, we introduce CL4D, a contrastive learning framework that continues pre-training the code generation models to enhance their representational capabilities. Follow Gao et al. (Gao et al. 2021), we employ hard negatives, semantically distinct code snippets that are close to the query in representation space, to improve training efficiency.
@@
To ensure rigorous evaluation, we compare pre-trained decoder-only and encoder-only models of comparable size on two representative understanding tasks: code search and clone detection. Experimental results indicate that our proposed CL4D effectively adapts decoder-only code generation models for code understanding tasks. After continued pre-training on a small code corpus, these models outperform pre-trained encoder-only models on most tasks. Furthermore, we observe that larger decoder-only models further improve code understanding performance.
@@
  • We investigate how pre-trained decoder-only code generation models can be effectively leveraged for code understanding tasks, bridging the gap between generation-oriented and understanding-oriented applications.
  • We propose CL4D, a contrastive learning–based continued pre-training approach that enhances the representation ability of decoder-only models, enabling competitive or superior performance to encoder-only models on code understanding benchmarks.
@@
  • Our findings reveal that larger decoder-only models benefit code understanding tasks and demonstrate the potential for unifying code understanding and code generation within a single model architecture.
Figure 2: Overview of CL4D.
Background and Motivation
In recent years, large-scale language models with a decoder-only architecture have achieved remarkable success in the field of code generation. Models such as Codex (Chen et al. 2021), CodeGen (Nijkamp et al. 2023), StarCoder (Li et al. 2023), and Code Llama (Rozière et al. 2023) have demonstrated impressive capabilities in generating syntactically correct and semantically meaningful code. These models are typically pre-trained on massive corpora of natural language and programming languages, allowing them to acquire extensive programming knowledge and multi-language fluency. With the emergence of instruction-tuned variants like WaveCoder (Yu et al. 2024) and CodeGemma (Zhao et al. 2024), decoder-only models continue to push the boundaries of generative performance.
Despite their superiority in generative tasks, our empirical observations reveal a surprising performance gap when these models are applied directly to code understanding tasks such as code search and code clone detection. As shown in Table 4, encoder-only models—such as GraphCodeBERT (Guo et al. 2021), UniXcoder (Guo et al. 2022) and CodeSage (Zhang et al. 2024)—which are explicitly designed for understanding source code, consistently outperform their decoder-only counterparts on benchmark code understanding datasets.
This performance gap persists even though decoder-only models possess several natural advantages. First, they have been exposed to significantly larger and more diverse code corpora during pre-training, often spanning multiple programming languages and domains. Second, these models are frequently trained with objectives aligned with next-token prediction, which implicitly teaches structural and semantic patterns in code. However, their lack of bi-directional attention mechanisms and their generative training objective may limit their capacity for fine-grained code comprehension.
One obvious solution is to pre-train a new encoder-only model at the same scale as the latest decoder-only models. However, the computational and financial costs of such an endeavor are prohibitively high, particularly when targeting billion-parameter models trained over trillions of tokens. This raises a critical research question: Can we enhance the code comprehension capabilities of existing decoder-only models without incurring the massive cost of training new models from scratch?
To this end, we propose a contrastive learning framework aimed at improving the code understanding performance of decoder-only models. Our approach leverages the inherent strengths of pre-trained decoder-only models while avoiding the prohibitive cost of full model retraining.
Methodology
@@
In this section, we explore obtaining encoding from decoder-only pre-trained code generation models to address code understanding tasks. The overview of CL4D is shown in Figure 2. We first introduce the data construction method and then introduce the training of CL4D, including the model architecture and contrastive learning.
Data Construction
@@
We train decoder-only Transformer models on six programming languages from The Stack dataset (Kocetkov et al. 2022): Python, Java, Go, PHP, JavaScript, and Ruby. To enhance the ability of model to differentiate between natural language and programming language, we use Tree-Sitter [^3] https://tree-sitter.github.io/tree-sitter/ to extract bimodal data from The Stack dataset for training, i.e., $(query,code)$ pairs. The query is the first line of the function docstring, which approximates the query of user. Additionally, we apply the filtering rules to improve the quality of the training corpus following CSN (Husain et al. 2019). Given that pre-trained code generation models have already learned a vast amount of code knowledge, we construct millions of data samples for continued pre-training.
Model Architecture
We investigated two methods for obtaining code representation from decoder-only models: 1) using the embedding of the last token, and 2) averaging the embeddings of all tokens. Encoder-only models like BERT typically use the embedding of the first token $[CLS]$ in the last hidden layer as the feature representation of the input code snippet, utilizing a bidirectional attention mechanism. By contrast, decoder-only models use a unidirectional attention mechanism, where only subsequent tokens can refer to previous information. Consequently, only the last token aggregates the information of the entire sample. In the first method, we use the embedding of the last token in the last hidden layer as the feature representation of the input code snippet. In the second method, we calculate the average of the embeddings of all input tokens in the last hidden layer as the feature representation.
Following previous work (Guo et al. 2021; Zhang et al. 2024), we adopt a dual-encoder architecture to continue pre-training the code encoding model. As depicted in Figure 2, this architecture consists of two Transformer decoder modules with shared weights to encode the input code snippet. In our experiments, we initialize the Transformer decoder module with the pre-trained code generation models.
Contrastive Learning
The traditional left-to-right training method for predicting the next token limits the representation capability of code generation models. To fully leverage the extensive code knowledge in the existing open-source code generation models and to avoid the high costs of retraining code understanding models, we employ contrastive learning to further fine-tune these models, thereby improving their representation abilities.
@@
Contrastive learning is the primary method for enhancing the representation ability of models (Zhang et al. 2023). Here, we employ contrastive learning tasks to train pre-trained code generation models, addressing the limited representation ability of decoder-only Transformer due to the unidirectional attention mechanism.
@@
Specifically, we randomly sample n samples from the training corpus to form a batch of data $B={(q_{1},c_{1}),...,(q_{n},c_{n})}$. For a $(q_{i},c_{i})$ pair, we take $c_{i}$ as the positive sample $c^{+}$ for $q_{i}$, and the other n1n-1 codes in the batch, excluding $c_{i}$, as negative samples for $q_{i}$. Additionally, following Gao et al. (Gao et al. 2021), we construct hard negative samples $c^{h}$ for $q_{i}$. Specifically, during the data processing stage, we utilize UniXcoder to select a code snippet from the entire codebase for each query in the training set, i.e., the code snippet that is close to the query in the representation space but have different semantic from the query. The loss function can be formulated as:
@@
  • RQ1: Can our proposed CL4D boost the performance of the decoder-only code generation models for the code understanding tasks?
  • RQ2: Which method is more suitable for a decoder-only model to extract code representation?
  • RQ3: To what extent does contrastive learning enhance the code representations of decoder-only models?
  • RQ4: How does CL4D perform in the zero-shot setting?
  • RQ5: Why does our approach work?
Understanding Tasks
We evaluate the code understanding capabilities of decoder-only Transformer on two primary tasks:
  • Code Search: This task aims to find the most semantically similar code snippet in a candidate codebase based on a given natural language query. We perform experiments on two datasets: CodeSearchNet (CSN) (Husain et al. 2019) and CoSQA (Huang et al. 2021). CSN is a multi-language code corpus collected from GitHub, including Python, Java, Go, PHP, JavaScript, and Ruby. CoSQA contains real user queries from Microsoft Bing search logs, with code snippets from CSN. We use Mean Reciprocal Rank (MRR) as the evaluation metric, which is widely adopted in previous studies (Guo et al. 2022; Zhang et al. 2024).
  • Clone Detection: This task aims to identify the most semantically similar code snippet in a candidate codebase based on a given code snippet as query. We conduct experiments on the POJ-104 dataset (Mou et al. 2016), designed to measure semantic similarity between code snippets. Mean Average Precision (MAP) is the evaluation metric for this task.
Baselines
We compare the code understanding capabilities of encoder-only models and decoder-only models of comparable sizes, and evaluate the extend to which the proposed CL4D method enhances the performance of decoder-only models.
@@
  • CodeBERT (Feng et al. 2020) (125M) is trained with masked language modeling and replaced token detection on six programming languages in CodeSearchNet.
  • GraphCodeBERT (Guo et al. 2021) (125M) is trained with additional code structure understanding tasks, including edge prediction and node alignment.
  • UniXcoder (Guo et al. 2022) (125M) is trained with multiple language modeling and contrastive learning tasks. It is an encoder-decoder architecture model, where only the encoder is used for code understanding tasks via a special token <encoder_only><encoder\_only>.
  • CodeSage (Zhang et al. 2024) (1.3B) is trained with masked language modeling and contrastive learning on the extensive code corpus from The Stack dataset.
  • CodeGPT (Lu et al. 2021) (125M) is trained on CodeSearchNet with the same structure and training objectives as GPT-2 (Radford et al. 2019).
  • CodeGen (Nijkamp et al. 2023) (350M) is a model for program analysis trained on extensive natural language and programming language data. We choose the 350M version to facilitate comparison with similarly sized encoder-only models.
  • SantaCoder (Allal et al. 2023) (1.1B) is trained with multi-query attention and fill-in-the-middle on Java, JavaScript, and Python subsets of The Stack, incorporating near-deduplication and comments-to-code filtering.
  • phi-1 (Gunasekar et al. 2023) (1.3B) is a code generation model trained using textbook data sourced from the internet and textbooks synthesized by GPT-3.5.
  • DeepSeek-Coder (Guo et al. 2024) (1.3B) is a model trained on a high-quality project-level code corpus.
Training details
During the pre-training phase, we utilize the tokenizer from each decoder-only model directly for tokenization. These models are trained using the AdamW optimizer, with a learning rate set to 2e-5. We train the decoder-only models for 2 epochs on the dataset, employing a batch size of 64. The experiments are conducted on 8 GPUs (80G A100). The longest pre-training duration is observed for phi-1, which takes approximately 3 days. In the fine-tuning phase, we adhere to the experimental settings of UniXcoder to minimize the impact of hyper-parameters on the fine-tuning results.
Evaluation
RQ1: Overall performance comparison
We evaluate the effectiveness of CL4D on two representative code understanding tasks: code search and clone detection. The experiment results are shown in Table 1, we can find that CL4D enables the decoder-only model to achieve state-of-the-art performance. Specifically, our method improves the performance of decoder-only models by approximately 2% compared to encoder-only models of the same size. Additionally, as the size of the decoder-only model increases, its performance on downstream code understanding tasks shows a corresponding improvement. Overall, our approach effectively enhances the code understanding capabilities of decoder-only models.
Method Code Search Clone Detection
CSN CoSQA POJ-104
Encoder-only CodeBERT (Feng et al. 2020) (125M) 70.18 65.7 83.79
GraphCodeBERT (Guo et al. 2021) (125M) 72.08 68.4 85.50
UniXcoder (Guo et al. 2022) (125M) 74.40 70.1 89.56
CodeSage (Zhang et al. 2024) (1.3B) 75.80 68.0 87.70
Decoder-only CodeGPT (Lu et al. 2021) (125M) + CL4D 70.20 69.0 87.96
CodeGen (Nijkamp et al. 2023) (350M) + CL4D 73.30 71.5 89.68
SantaCoder (Allal et al. 2023) (1.1B) + CL4D 74.98 72.2 83.98
phi-1 (Gunasekar et al. 2023) (1.3B) + CL4D 75.18 72.8 92.72
DeepSeek-Coder (Guo et al. 2024) (1.3B) + CL4D 77.57 71.9 89.71
Table 1: Comparison of the performance of encoder-only models and decoder-only models on code understanding tasks.
RQ2: Code representation extraction method study for decoder-only models
@@
To identify the most effective strategy for extracting code representations from decoder-only Transformer models, we investigated two approaches: decoder-only last token (Last) and decoder-only average (Avg). In addition, we examined how different padding strategies influence the quality and effectiveness of the resulting code representations.